Please help me solve this Question :
A certain amount of alcohol with a concentration of 90%
was mixed with an amount of alcohol smaller by 20 liters
and having a concentration of 30%. Afterward, 15 liters of
water evaporated from the mixture and the concentration of
the alcohol was 72%. How many liters of alcohol of 72%
concentration were formed?

Answer by lwsshak3(11628)

A certain amount of alcohol with a concentration of 90%
was mixed with an amount of alcohol smaller by 20 liters
and having a concentration of 30%. Afterward, 15 liters of
water evaporated from the mixture and the concentration of
the alcohol was 72%. How many liters of alcohol of 72%
concentration were formed?
**
let x=liters of 90% alcohol
x-20=liters of 30% alcohol
15=liters evaporated
(x+(x-20)-15)=(2x-35) liters of 72% concentration
..
.9x+.3(x-20)=.72(2x-35)
.9x+.3x-6=.72(2x-35)=1.44x-25.2
.24x=19.2
x=19.2/.24
x=80
2x-35=125
ans:
liters of alcohol of 72% concentration were formed=125 liters